The Everton College store paid $1700 for an order of 48 calculators. The store paid $8 for each scientific calculator. the others, all graphing calculators, cost the store $55 each. How many of each type of calculator was ordered?
x +y = 48
8 + 55 =1700 - now I get confused, can someone direct me?

Answer by rfer(16322)

s+g=48 g=48-s
--------------
8s+55g=1700
8s+55(48-s)=1700
8s+2640-55s=1700
-47s=-940
s=20 scientific
---------------
8*20+55g=1700
160+55g=1700
55g=1540
g=28 graphing